# Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges

### v1

@@ 5.1. Image and Video Captioning
VSG models need to understand the temporal progression of events, causal relationships, and the ability to construct cohesive and engaging stories. Therefore, they build upon the foundational principles of image and video captioning, incorporating both the interpretation of visual data and the generation of coherent, contextually appropriate narratives.
@@
The tasks of image and video captioning consist in generating descriptive, human-readable sentences that accurately represent the content of an image or video, identifying the main components in a visual scene and their relationships, alongside the ability to express these details in natural language. Initial efforts in image and video captioning relied on template-based and rule-based approaches. Current deep learning models are capable of end-to-end training, resulting in a substantial enhancement in the quality of the generated captions.
@@
One of the most notable contributions in this domain was the application of encoder-decoder architectures, known as the ‘show and tell’ approach (Vinyals et al., 2014; Yao et al., 2015; Anderson et al., 2017). This technique employs Convolutional Neural Networks (CNNs) as encoders to extract visual features. Specifically, 2D-CNNs are employed for image processing where the model interprets two dimensions: height and width. On the other hand, for video processing, 3D-CNNs are typically used where an additional dimension, time or depth, is taken into account to perceive the temporal dynamics in video sequences. Then, RNNs or their improved variants, LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Chung et al., 2014), known for their ability to effectively model sequential data, are commonly used as decoders in image and video captioning tasks, enabling the generation of coherent and contextually meaningful textual captions and descriptions from the visual features extracted by the CNNs. This combination of CNNs and RNNs effectively bridges the gap between vision and language.
@@
In the realm of image captioning, transformers have been used as encoders to process visual features extracted from images. Leveraging the self-attention mechanism inherent in transformers, these models are able to capture relationships between visual elements, attending to relevant regions in the image, and highlighting visual cues and their interdependencies (Olimov et al., 2021; Radford et al., 2021; Liu et al., 2021a; Yu et al., 2019). Transformer models also demonstrate the ability to generate precise and elaborate textual descriptions, significantly improving the overall quality of image captions.
@@
Similarly, transformers have made contributions to video captioning, addressing the challenges associated with understanding and describing temporal dynamics in videos. By treating video frames as a sequential series of images, transformers can leverage their attention mechanism to capture long-range dependencies across frames (Liu et al., 2021b; Lin et al., 2021) This enables the models to map the temporal progression of events, infer causal relationships, and grasp the underlying narrative structure within the video.
@@
The integration of transformers in image and video captioning has yielded improvements that surpass the performance of traditional encoder-decoder architectures (Liu et al., 2021a, b). Transformers not only generate semantically accurate captions but also exhibit enhanced coherence and contextual relevance. Their attention mechanism allows for a incorporation of local and global context, enabling a more holistic understanding of the visual content and facilitating caption generation.
@@
The use of transformers in image and video captioning holds great promise, but there are challenges to be addressed. Efficient fusion of visual features with textual information (Radford et al., 2021), handling long-term dependencies in video sequences (Liu et al., 2021b), and ensuring temporal coherence in caption generation for longer videos are among the ongoing research topics.
5.2. Visual Question Answering
@@
Visual Question Answering (VQA) combines computer vision and natural language processing to enable the extraction of information from visual content, identifying objects, scenes, actions, and comprehending context, and provide meaningful answers to questions about that content. Both VQA and VSG involve understanding visual content and generating meaningful responses and the developments in VQA can be directly applied in VSG tasks. Through the use of VQA techniques, VSG models can gain valuable insights into the content and context of images and videos, improving the overall quality of their visual narratives.
@@
Encoder-classifier deep-learning VQA models encode visual and textual inputs and select an answer from a predefined set using a classification process. Image encoders, typically pretrained CNNs, extract visual features from the input images. Question encoders, often RNNs, processes the input questions, converting them into fixed-length latent representations. Fusion techniques combine the visual and textual features, creating a fused representation that captures their interactions that is then classified using fully connected layers, mapping them to possible answer classes. Traditional VQA models are trained using supervised learning with large-scale datasets, where image-question-answer triplets are provided as training examples. The parameters are optimized to minimize the loss between predicted answers and ground truth answers during training. Some examples are BUTD (Anderson et al., 2017), that uses an attention-based model that combines bottom-up and top-down mechanisms to improve performance in image captioning and visual question answering tasks, and “Show, Ask, Attend, and Answer” (Kazemi and Elqursh, 2017), that combines LSTMs with multiple attention layers.
@@
The transition from encoder-classifier methods to transformer methods in VQA was driven by the transformer’s ability to capture long-range dependencies and global contextual information. Transformers offered the potential to more efficiently capture the interactions between visual and textual modalities, enabling a more comprehensive understanding of the relationship between images and questions. In these VQA models, images and questions are encoded using transformer encoders and self-attention mechanisms capture the dependencies between words and the visual features, allowing the model to focus on relevant image regions while processing the question. The fused encoded representations are then fed into classification layers to generate answers (Bao et al., 2022; Wang et al., 2023a, 2022, b). Recent advances in VQA also explore new models that approach the task as a generation problem rather than a classification one, allowing for a more flexible and diverse range of answers (Chen et al., 2022).
6. Visual Story Generation
The current state-of-the-art and baseline methods employed in VSG have evolved from their origins in textual story generation. These methods have served as foundations for the development of more sophisticated techniques and continue to provide a comparative baseline for new models and approaches.
6.1. Visual Storytelling with Convolutional and Recurrent Neural Networks
@@
Huang et al. (Huang et al., 2016) proposed a dataset of sequential vision-to-language mappings, SIND v.1 (later renamed to VIST), with 81,743 images in 20,211 sequences, with both descriptive (captions) and story language, and aiming at facilitating research, from understanding individual images to comprehending image sequences that narrate events over time.
@@
This work also proposed models for generating stories based on image sequences, using a sequence-to-sequence RNN approach, which extends the image captioning technique. It extracts features from each image using a pretrain CNN called AlexNET (Krizhevsky et al., 2012). It then encodes image sequences by using a GRU over the features of each individual image. The story decoder model consists of another GRU and generates narratives word by word. Four different decoding methods, shown in Table 3, were considered: beam search, greedy search, greedy search with removal of duplicates, and greedy search with removal of duplicates and incorporation of grounded words.
Decoding Method METEOR
@@
Table 3. Results of each decoding approach using evaluated using METEOR scores on the VIST (Huang et al., 2016) dataset.
@@
Figure 7. Sequence of images used to generate stories using different decoding methods. Beam search: This is a picture of a family. This is a picture of a cake. This is a picture of a dog. This is a picture of a beach. This is a picture of a beach. Greedy: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The dog was enjoying the water. The dog was happy to be in the water. Greedy - duplicates: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The kids were playing in the water. The boat was a little too much to drink. Greedy - duplicates + grounded words: The family got together for a cookout. They had a lot of delicious food. The dog was happy to be there. They had a great time on the beach. They even had a swim in the water. Adapted from (Huang et al., 2016).
6.2. Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
Park et al. (Park and Kim, 2015) proposed Coherent Recurrent Convolutional Network (CRCN), integrating various neural network components to bridge the gap between visual content and textual narrative: VGGNet (Simonyan and Zisserman, 2015) as image encoder and Bidirectional Recurrent Networks (BRNN) to capture contextual information from both the forward and backward directions within a given text sequence. CRCN also introduces a local coherence model that focuses on maintaining coherence and context within the generated sentence sequences. It also utilizes parse trees to detect grammatical roles and structural attributes, ensuring that the generated sentences integrate with the broader narrative context. During training, correctly aligned image-sentence pairs are given higher scores than misaligned pairs.
The evaluation of CRCN considered several automatic metrics and several baseline works, including one that excludes the entity coherence model. Table 4 shows the results for the NY City (Park and Kim, 2015) and Disneyland (Park and Kim, 2015) datasets. Both datasets were introduced by Park et al. (Park and Kim, 2015) and consist of image sequences with the corresponding blog posts. In these datasets, results were superior to all baselines. The exception, where it was a close second, was with the baseline without entity coherence and in only certain metrics. The work was further validated by user studies conducted via Amazon Mechanical Turk (AMT), showing that human evaluators prefer CRCN-generated sequences over several baselines.
@@ 6.2. Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
Table 4. Performance Metrics for CRCN on New York City and Disneyland Datasets. Bold values represent state-of-the-art results at the time. B stands for BLEU, C for CIDEr, M for METEOR and R for Recall (Park and Kim, 2015).
@@
Figure 8. Sequence of images used by CRCN a story. (1) One of the hallway arches inside of the library (2) As we walked through the library I noticed an exhibit called lunch hour nyc it captured my attention as I had also taken a tour of nyc food carts during my trip (3) Here is the top of the Chrysler building everyone’s favorite skyscraper in new york. (4) After leaving the nypl we walked along 42nd st. (5) We walked down fifth avenue from rockefeller centre checking out the windows in saks the designer stores and eventually making our way to the impressive new york public library. Adapted from (Park and Kim, 2015).
Besides, Park et al. (Park and Kim, 2015) which is the first approach to use a bidirectional RNN for visual storytelling, to the best of our knowledge, XE (Wang et al., 2018) and AREL (Wang et al., 2018) have also proposed models for visual storytelling that use bidirectional RNNs on the VIST dataset.
6.3. Transitional Adaptation of Pretrained Models for Visual Storytelling
Yu et al. (Yu et al., 2021) proposed Transitional Adaptation of Pretrained Model (TAPM), aimed at refining the generation of textual descriptions for visual content, particularly in the context of VSG tasks. It aims at bridging the gap between pretrained language models and visual encoders. In contrast with previous models for vision-to-language generation tasks, which typically pretrain a visual encoder and a language generator separately and then jointly fine-tune them for the target task, TAPM proposes a transitional adaptation task to harmonize the visual encoder and language model for downstream tasks like VSG, mitigating discord between visual specificity and language fluency arising from separate training on large corpora of visual and text data. TAPM introduces an approach that adapts multimodal modules through a simpler alignment task focusing solely on visual inputs, eliminating the need for text labels.
@@ 6.3. Transitional Adaptation of Pretrained Models for Visual Storytelling
The loss function divides sequential coherence into three components: past, current, and future matching losses. The past matching loss uses a Fully Connected (FC) layer $f_{p}$ to project the text representation $b_{s_{i}}$ of video $i$, drawing it nearer to the visual representation $v_{b_{i-1}}$ of the preceding video $i-1$ and distancing it from those of non-sequential videos. The future matching loss projects $b_{s_{i}}$ via a distinct FC layer $f_{f}$, aligns with the subsequent visual representation $v_{b_{i+1}}$. The current matching loss then aligns the current visual representation $v_{b_{i}}$ with $b_{s_{i}}$ through another FC layer $f_{c}$. These components are unified by the FC layer projections in their respective visual spaces, pulling the embeddings of correct matches closer and pushing incorrect matches further apart. Margin ranking losses are utilized to implement this concept, contrasting correct matches with incorrect ones. The final sequential coherence loss for a given video $i$ is formulated as shown in Eq. 15, where $\cos$ is the cosine similarity, and $j$ represents indices of incorrect matches.
(15) $L_{i}=$ $\sum_{j\neq i-1}\max(0,1+\cos(v_{b_{j}},f_{p}(b_{s_{i}}))-\cos(v_{b_{i-1}},f_{p}(b_{s_{i}})))$
@@
Fig. 9(a) shows results for TAPM against selected baselines on the LSMDC 2019 and VIST datasets. Fig. 9(b) and 9(c) show human evaluation results in which TAPM surpasses adversarial baselines on LSMDC 2019. On VIST it surpasses the XE (Wang et al., 2018) and AREL (Wang et al., 2018) baselines in relevance, expressiveness, and concreteness. These results highlight the strengths in word choice and contextual accuracy, showcasing its ability to capture causal relationships between images. However, the score is still far from human performance, indicating that there is still room for improvement.

### v2

@@ 4.1 Image and Video Captioning
VSG models need to understand the temporal progression of events, causal relationships, and the ability to construct cohesive and engaging stories. VSG can also speculate on the motivations and emotions of characters since such interpretations do not contradict the visual content. Therefore, they build upon the foundational principles of image and video captioning, incorporating both the interpretation of visual data and the generation of coherent, contextually appropriate narratives.
@@
The tasks of image and video captioning consist in generating descriptive, human-readable sentences that represent the content of an image or video, identifying the main components in a visual scene and their relationships, alongside the ability to express these details in natural language.
@@
One notable contribution in this domain was the application of encoder-decoder architectures, known as the ‘show and tell’ approach Vinyals et al. (2014); Yao et al. (2015); Anderson et al. (2017). This technique employs Convolutional Neural Networks (CNNs) as encoders to extract visual features. Specifically, 2D-CNNs are employed for image processing where the model interprets two dimensions: height and width. On the other hand, for video processing, 3D-CNNs are typically used where an additional dimension, time or depth, is taken into account to perceive the temporal dynamics in video sequences. Then, Recurrent Neural Networks (RNNs) or their improved variants, Long Short-Term Memory networks (LSTMs) Hochreiter and Schmidhuber (1997) and Gated Recurrent Units (GRUs) Chung et al. (2014), known for their ability to effectively model sequential data, are commonly used as decoders in image and video captioning tasks, enabling the generation of coherent and contextually meaningful textual captions and descriptions from the visual features extracted by the CNNs. This combination of CNNs and RNNs effectively bridges the gap between vision and language.
@@
In the realm of image captioning, Transformers Vaswani et al. (2017) have been used as encoders to process visual features extracted from images. Leveraging the self-attention mechanism inherent in Transformers, these models are able to capture relationships between visual elements, attending to relevant regions in the image, and highlighting visual cues and their interdependencies Olimov et al. (2021); Radford et al. (2021); Liu et al. (2021); Yu et al. (2019). Transformer models also demonstrate the ability to generate precise and elaborate textual descriptions, significantly improving the overall quality of image captions.
@@
Similarly, Transformers have made contributions to video captioning, addressing the challenges associated with understanding and describing temporal dynamics in videos. By treating video frames as a sequential series of images, Transformers can leverage their attention mechanism to capture long-range dependencies across frames Liu et al. (2021); Lin et al. (2021) This enables the models to map the temporal progression of events, infer causal relationships, and grasp the underlying narrative structure within the video.
@@
The integration of Transformers in image and video captioning has yielded improvements that surpass the performance of traditional encoder-decoder architectures Liu et al. (2021, 2021). Transformers not only generate semantically accurate captions but also exhibit enhanced coherence and contextual relevance. Their attention mechanism allows for a incorporation of local and global context, enabling a more holistic understanding of the visual content and facilitating caption generation.
@@
The use of Transformers in image and video captioning holds great promise, but there are challenges to be addressed. Efficient fusion of visual features with textual information Radford et al. (2021), handling long-term dependencies in video sequences Liu et al. (2021), and ensuring temporal coherence in caption generation for longer videos are among the ongoing research topics.
4.2 Visual Question Answering
@@
Visual Question Answering (VQA) combines computer vision and NLP to enable the extraction of information from visual content, identifying objects, scenes, actions, and comprehending context, and provide meaningful answers to questions about that content. Both VQA and VSG involve understanding visual content and generating meaningful responses and the developments in VQA can be directly applied in VSG tasks. Through the use of VQA techniques, VSG models can gain valuable insights into the content and context of images and videos, improving the overall quality of their visual narratives.
@@
Encoder-classifier deep-learning VQA models encode visual and textual inputs and select an answer from a predefined set using a classification process. Image encoders, typically pretrained CNNs, extract visual features from the input images. Question encoders, often RNNs, processes the input questions, converting them into fixed-length latent representations. Fusion techniques combine the visual and textual features, creating a fused representation that captures their interactions that is then classified using fully connected layers, mapping them to possible answer classes. Traditional VQA models are trained using supervised learning with large-scale datasets, where image-question-answer triplets are provided as training examples. The parameters are optimized to minimize the loss between predicted answers and ground truth answers during training. Some examples are BUTD Anderson et al. (2017), that uses an attention-based model that combines bottom-up and top-down mechanisms to improve performance in image captioning and visual question answering tasks, and “Show, Ask, Attend, and Answer” Kazemi and Elqursh (2017), that combines LSTMs with multiple attention layers.
@@
The transition from encoder-classifier methods to Transformer methods in VQA was driven by the Transformer’s ability to capture long-range dependencies and global contextual information. Transformers offered the potential to more efficiently capture the interactions between visual and textual modalities, enabling a more comprehensive understanding of the relationship between images and questions. In these VQA models, images and questions are encoded using Transformer encoders and self-attention mechanisms capture the dependencies between words and the visual features, allowing the model to focus on relevant image regions while processing the question. The fused encoded representations are then fed into classification layers to generate answers Bao et al. (2022); Wang et al. (2023, 2022, 2023). Recent advances in VQA also explore new models that approach the task as a generation problem rather than a classification one, allowing for a more flexible and diverse range of answers Chen et al. (2022).
5 Visual Story Generation
VSG combines computer vision and NLP to create coherent narratives from sequences of images or videos. This task extends beyond simple image captioning by requiring models to understand temporal relationships, infer causal connections between visual scenes, and generate engaging stories that capture the underlying narrative flow. Unlike video captioning, which focuses on objective description of visual content, VSG can also speculate on the motivations and emotions of characters since such interpretations do not contradict the visual content.
The field has evolved from early encoder-decoder architectures using CNN and RNN to Transformer-based models that leverage large-scale pretraining. These approaches must address challenges including visual understanding, temporal coherence, and narrative creativity while maintaining grounding in the provided visual content. The following subsections present the key methodologies and their contributions to advancing VSG capabilities.
5.1 Visual Storytelling with Convolutional and Recurrent Neural Networks
@@
Huang et al. Huang et al. (2016) proposed a dataset of sequential vision-to-language mappings, SIND v.1 (later renamed to VIST), with 81,743 images in 20,211 sequences, with both descriptive (captions) and story language, and aiming at facilitating research, from understanding individual images to comprehending image sequences that narrate events over time.
@@
This work also proposed models for generating stories based on image sequences, using a sequence-to-sequence RNN approach, which extended the image captioning technique. It extracts features from each image using a pretrain CNN called AlexNET Krizhevsky et al. (2012). It then encodes image sequences by using a GRU over the features of each individual image. The story decoder model consists of another GRU and generates narratives word by word. Four different decoding methods, shown in Table 3, were considered: beam search, greedy search, greedy search with removal of duplicates, and greedy search with removal of duplicates and incorporation of grounded words.
Method METEOR Score
@@
Table 3: Results of each decoding approach using evaluated using METEOR scores on the VIST Huang et al. (2016) dataset. Red indicates the lowest score, green the highest.
@@
Figure 1: Sequence of images used to generate stories using different decoding methods. Beam search: This is a picture of a family. This is a picture of a cake. This is a picture of a dog. This is a picture of a beach. This is a picture of a beach. Greedy: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The dog was enjoying the water. The dog was happy to be in the water. Greedy - duplicates: The family gathered together for a meal. The food was delicious. The dog was excited to be there. The kids were playing in the water. The boat was a little too much to drink. Greedy - duplicates + grounded words: The family got together for a cookout. They had a lot of delicious food. The dog was happy to be there. They had a great time on the beach. They even had a swim in the water. Adapted from Huang et al. (2016).
5.2 Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
Park et al. Park and Kim (2015) proposed Coherent Recurrent Convolutional Network (CRCN), integrating various neural network components to bridge the gap between visual content and textual narrative: VGGNet Simonyan and Zisserman (2015) as image encoder and Bidirectional Recurrent Networks (BRNN) to capture contextual information from both the forward and backward directions within a given text sequence. CRCN also introduces a local coherence model that focuses on maintaining coherence and context within the generated sentence sequences. It also utilizes parse trees to detect grammatical roles and structural attributes, ensuring that the generated sentences integrate with the broader narrative context. During training, correctly aligned image-sentence pairs are given higher scores than misaligned pairs.
The evaluation of CRCN considered several automatic metrics and several baseline works, including one that excludes the entity coherence model. Table 4 shows the results for the NY City Park and Kim (2015) and Disneyland Park and Kim (2015) datasets. Both datasets were introduced by Park et al. Park and Kim (2015) and consist of image sequences with the corresponding blog posts. In these datasets, results were superior to all baselines. The exception, where it was a close second, was with the baseline without entity coherence and in only certain metrics. The work was further validated by user studies conducted via Amazon Mechanical Turk (AMT), showing that human evaluators prefer CRCN-generated sequences over several baselines.
@@ 5.2 Visual Storytelling with Convolutional and Bidirectional Recurrent Networks Networks
Table 4: Performance metrics of CRCN on New York City and Disneyland Datasets. Green values represent state-of-the-art results at the time. B stands for BLEU, C for CIDEr, M for METEOR and R for Recall Park and Kim (2015).
@@
Figure 2: Sequence of images used by CRCN in a story. (1) One of the hallway arches inside of the library (2) As we walked through the library I noticed an exhibit called lunch hour nyc it captured my attention as I had also taken a tour of nyc food carts during my trip (3) Here is the top of the Chrysler building everyone’s favorite skyscraper in new york. (4) After leaving the nypl we walked along 42nd st. (5) We walked down fifth avenue from rockefeller centre checking out the windows in saks the designer stores and eventually making our way to the impressive new york public library. Adapted from Park and Kim (2015).
In addition to the work by Park et al. Park and Kim (2015), which was the first to use a bidirectional RNN for visual storytelling, other works have also employed bidirectional RNNs architectures. XE Wang et al. (2018) and AREL Wang et al. (2018) have proposed models for visual storytelling that use bidirectional RNNs on the VIST dataset.
5.3 Transitional Adaptation of Pretrained Models for Visual Storytelling
Yu et al. Yu et al. (2021) proposed Transitional Adaptation of Pretrained Model (TAPM), aimed at refining the generation of textual descriptions for visual content, particularly in the context of VSG tasks. It aims at bridging the gap between pretrained language models and visual encoders. In contrast with previous models for vision-to-language generation tasks, which typically pretrain a visual encoder and a language generator separately and then jointly fine-tune them for the target task, TAPM proposes a transitional adaptation task to harmonize the visual encoder and language model for downstream tasks like VSG, mitigating discord between visual specificity and language fluency arising from separate training on large corpora of visual and text data. TAPM introduces an approach that adapts multimodal modules through a simpler alignment task focusing solely on visual inputs, eliminating the need for text labels.
@@ 5.3 Transitional Adaptation of Pretrained Models for Visual Storytelling
The loss function divides sequential coherence into three components: past, current, and future matching losses. The past matching loss uses a Fully Connected (FC) layer $f_{p}$ to project the text representation $b_{s_{i}}$ of video $i$, drawing it nearer to the visual representation $v_{b_{i-1}}$ of the preceding video i1i-1 and distancing it from those of non-sequential videos. The future matching loss projects $b_{s_{i}}$ via a distinct FC layer $f_{f}$, aligns with the subsequent visual representation $v_{b_{i+1}}$. The current matching loss then aligns the current visual representation $v_{b_{i}}$ with $b_{s_{i}}$ through another FC layer $f_{c}$. These components are unified by the FC layer projections in their respective visual spaces, pulling the embeddings of correct matches closer and pushing incorrect matches further apart. Margin ranking losses are utilized to implement this concept, contrasting correct matches with incorrect ones. The final sequential coherence loss for a given video $i$ is formulated as shown in Eq. 1, where $\cos$ is the cosine similarity, and $j$ represents indices of incorrect matches.
$L_{i}=$ $\sum_{j\neq i-1}\max(0,1+\cos(v_{b_{j}},f_{p}(b_{s_{i}}))-\cos(v_{b_{i-1}},f_{p}(b_{s_{i}})))$ (1)
@@
Table. 6 and 6 shows results for TAPM against selected baselines on the LSMDC 2019 and VIST datasets. Table. 7 show human evaluation results in which TAPM surpasses adversarial baselines on LSMDC 2019. On VIST it surpasses the XE Wang et al. (2018) and AREL Wang et al. (2018) baselines in relevance, expressiveness, and concreteness as shown in Table 8. These results highlight the strengths in word choice and contextual accuracy, showcasing its ability to capture causal relationships between images. However, the score is still far from human performance, indicating that there is still room for improvement.
Models C M R
Huang et al. Huang et al. (2016) - 31.4 -
AREL Wang et al. (2018) 9.4 35.0 29.5
StoryAnchor Modi and Parde (2019) 9.9 35.5 30.0
HSRL Huang et al. (2018) 10.7 35.2 30.8
INet Jung et al. (2020) 10.0 35.6 29.7
TAPM 13.8 37.2 33.1
Table 5: Quantitative results on the VIST test set. C stands for CIDEr, M for METEOR, and R for ROUGE-L. Green indicates the highest score, red the lowest.